A solenoid carries a current of 9.2A and has a coil density of 977turns/cm. What is the magnitude of the magnetic field near the middle of the solenoid? (mT, three significant figures)
Answer #1


the magnetic field near the middle of the solenoid

B = uo*n*I


n = number of turns per unit length = 977 turns/ cm = 97700 turns / m


I = current through the solenoid


uo = magnetic permeability of the free space = 4*pi*10^-7


B = 4*3.14*10^-7*97700*9.2

B = 1128.9 mT